OpenStudy (r_josh10):

Variable y varies directly with variable x, and y = 4 when x = 7. What is the value of x when y = 12? Is it 21? Im not Sure

OpenStudy (anonymous):

y = kx Plug in what we know: 4 = 7k Divide 7 to both sides: k = 4/7 So this is the constant, plug in what we know: y = kx It says when 'y = 12' 12 = 4/7x Divide 4/7 to both sides or multiply by its reciprocal, 7/4. 7/4 * 12 = 21 x = 21 Yes, 21 is correct!

OpenStudy (anonymous):

@r_josh10

OpenStudy (r_josh10):

Thank you Again You are the most helpful

OpenStudy (r_josh10):

@iGreen